OpenStudy (goodday):

3. Find the product of 5×3/5 .

OpenStudy (usukidoll):



OpenStudy (usukidoll):

is that your question

OpenStudy (goodday):

yes

OpenStudy (usukidoll):

what's 5 x 3?

OpenStudy (goodday):

what is the answer

OpenStudy (usukidoll):

I can't give direct answers

pooja195 (pooja195):

Hi welcome to openstudy :) Let's work through the problem rather than trying to get answers only. @UsukiDoll is helping you out. Listen to her.

OpenStudy (goodday):

ok

TheSmartOne (thesmartone):

You can quickly solve your question by cancelling the 5 since we have it in the numerator and denominator :P

OpenStudy (usukidoll):

O_O good point

OpenStudy (shamim):

plz tell me \[1 \times 2 =?\]

OpenStudy (shamim):

@goodday

OpenStudy (shamim):

u can also try \[4 \times \frac{ 3 }{ 4 } =?\]

OpenStudy (photon336):

in case you're wondering: 5 is actually 5/1 you multiply across the top and bottom \[(\frac{ 5 }{ 1 })*(\frac{ 3 }{ 5 }) = \frac{ }{ } \]

TheSmartOne (thesmartone):